OpenStudy (anonymous):

let f be the function satisfying f'(x)= xsqrtf(x), for all real numbers , where f(3)=25. Find f"(3). i need help with this :/

jimthompson5910 (jim_thompson5910):

\[\Large f \ ' (x) = x*\sqrt{f(x)}\] right?

OpenStudy (anonymous):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the value of f ' (3) ? Are you able to compute this?

OpenStudy (anonymous):

3*sqrt f(3)

jimthompson5910 (jim_thompson5910):

\[\Large f \ ' (x) = x*\sqrt{f(x)}\] \[\Large f \ ' (3) = 3*\sqrt{f(3)}\] yep, keep going

jimthompson5910 (jim_thompson5910):

since we're given f(3) = 25, you can replace f(3) with 25 and simplify further

OpenStudy (anonymous):

okay so f'(3)=15?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

that will be used when computing f '' (3)

OpenStudy (anonymous):

okay what is the next step

jimthompson5910 (jim_thompson5910):

\[\Large f \ ' (x) = x*\sqrt{f(x)}\] \[\Large f \ '' (x) = \sqrt{f(x)}+x*\frac{1}{2\sqrt{f(x)}}*f \ ' (x)\] \[\Large f \ '' (3) = \sqrt{f(3)}+3*\frac{1}{2\sqrt{f(3)}}*f \ ' (3)\] \[\Large f \ '' (3) = ???\] I used the product rule to derive x*sqrt(f(x)) on step 2

jimthompson5910 (jim_thompson5910):

also, the chain rule is being applied on step 2

OpenStudy (anonymous):

so i got 5+3(1/10)(15)

jimthompson5910 (jim_thompson5910):

looks good what do you get when you simplify that into one fraction?

OpenStudy (anonymous):

19/2

jimthompson5910 (jim_thompson5910):

I'm getting the same thing

OpenStudy (anonymous):

okay well thank u :)

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

are you sure 5+ (3/10)*15 =5+9/2= 9(1/2)